OpenStudy (anonymous):

p/5-p w/ a underline 4p/2p+1

OpenStudy (anonymous):

\[\left(\frac{p}{5}-\frac{p w}{a}\right)/\left(\frac{4p}{2p+1}\right)=\frac{(2 p+1) (a-5 w)}{20 a} \]